OpenStudy (anonymous):

Solve for x in the proportion x over the quantity 2 times x plus 5 equals negative 5 over x. x = 5 x = 15 x = −25 x = −5

jimthompson5910 (jim_thompson5910):

x/(2x+5) = -5/x x*x = -5(2x+5) x^2 = -10x - 25 x^2 + 10x + 25 = 0 I'll let you finish up

OpenStudy (anonymous):

\[\frac{ x }{ 2x + 5} = \frac{ -5 }{ x }\]

jimthompson5910 (jim_thompson5910):

do you know how to take it from here?

OpenStudy (anonymous):

No :/

jimthompson5910 (jim_thompson5910):

x^2 + 10x + 25 = 0 (x+5)(x+5) = 0 ... factor the left side x+5 = 0 or x+5 = 0 x = -5 or x = -5 so the only solution is x = -5